OpenStudy (anonymous):

One beam of electrons moves at right angles to a magnetic field. The force on these electrons is 4.9 × 10-14 newtons. A second beam travels at the same speed, but at a 30° angle with the magnetic field. What force is on these electrons?Select one of the options below as your answer: A. (4.9 × 10-14 newtons) · tan(30°) B. (4.9 × 10-14 newtons) · sin(30°) C. (4.9 × 10-14 newtons) · cos(30°) D. (4.9 × 10-14 newtons) · arctan(30°) E. (4.9 × 10-14 newtons) · arccos(30°)

OpenStudy (souvik):

there is two situation... 1)F=qvBsin(theta) in this situation F=4.9*10^-14 N theta =90 degree.....so sin(theta)=1 q(magnitude)=1.6*10^-19 C find B

OpenStudy (souvik):

in the second situation you have to find F you know q,theta,B ...

OpenStudy (anonymous):

sin30?

OpenStudy (souvik):

sin30=0.5

OpenStudy (anonymous):

well i guess ill go with that

OpenStudy (souvik):

well...you guessed it right...

OpenStudy (anonymous):

kool